Eight 7.0-W Christmas tree lights are connected in series to each other and to a 110-V source. (a) Show the illustration (b) What is the resistance of each bulb?


1
Expert's answer
2022-04-06T10:13:36-0400

Part (a)



(B)

Total power

"P=7\times8=56W"

Now current

"I=\frac{56}{110}=0.509A"

"R_1=\frac{P}{I^2}=\frac{7}{0.509^2}=27.01\Omega"

Similarly

"R_2=\frac{P}{I^2}=\frac{7}{0.509^2}=27.01\Omega"


"R_3=R_4=R_5=R_6=R_7=R_8=27.01\Omega"